I am going to choose 6 students out of the 15 who currently have A's to be mentors for the students who have C's or lower. How many ways can a group of 6 be selected?
Answer by vleith(2983)

Use a combination. See http://www.mathagonyaunt.co.uk/STATISTICS/ESP/Perms_combs.html
15P6 = 15!/((15-6)! * 6!)
= (15*14*13*12*11*10)/ (6*5*4*3*2*1)
= (14*13*11*10)/4
= 7*13*11*5
= 5005